If a=0.65 and b=1.28 and c= -5/2,evaluate -a(-b)+c.I don't get what to do at all,my answer I get is -5.83 and that's wrong as I was told can you help me? And I have two more problems and that's it.The next one is solve the equation for x, 6x minus 18=4x+8 and I don't know where to begin.And the last one is solve the equation for n,-55 minus 4n=13,there is another problem that I don't know where to begin can you please help me, thank you very much.
Found 2 solutions by stanbon, chitra:
Answer by stanbon(48535)

If a=0.65 and b=1.28 and c= -5/2,evaluate -a(-b)+c.
-0.65(-1.28)+(-5/2)
=0.832-(5/2)
=0.832-2.5
=-1.668
------------------------
The next one is solve the equation for x, 6x minus 18=4x+8
6x-18=4x+8
Subtract 4x from both sides to get:
2x-18=8
Add 18 to both sides to get:
2x=26
Divide both sides by 2 to get:
x=13
-------------
And the last one is solve the equation for n,-55 minus 4n=13
-55-4n=13
Add 55 to both sides to get:
-4n=42
Divide both sides by -4 to get:
n=-10.5
--------
Cheers,
Stan H.

Answer by chitra(359)

The first problem is:

-a(-b) + c given a = 0.65, b = 1.28 and c = -5/2


-0.65(-1.28) + (-5/2)

-0.65(-1.28) - (5/2) [Multiply the first two terms and product of two negatives is a positive]

0.832 - 2.5 [Subtract each other]

= -1.668 [negative since, 2.5 is greater and has a negative sign]

--------------------------------------------------------------------------

Solve for x:

6x - 18 = 4x + 8

Grouping the like terms, we get:

That is x on one side and numbers on the other side.

Adding 18 - 4x on both sides, we get:


6x - 4x = 8 + 18

2x = 26

x = 26/2

x = 13

-------------------------------------------------------------------
Solve for n:

-55 - 4n = 13

Adding 4n - 13 on both sides of the above equation, we get:

-55 - 13 = 4n

-68 = 4n

hence, we find out the value of "n"

n = -68/4

Therefore, n = -17

-------------------------------------------------------------------------
Hence, the solutions.

If you have any more questions and queries you can mail it to my ID. My ID is
tutyfruty9@yahoo.com